OpenStudy (anonymous):

Challenge problem , show that for any n positive even integer \(\large \gcd(3434,(\frac{10^{6n-6}-1}{99}) )= 1 \)

ganeshie8 (ganeshie8):

99 is below the entire thing right ?

OpenStudy (anonymous):

yes that gives u 99999999.....9/99=1010101....01

ganeshie8 (ganeshie8):

look at the expression in question it is wrong

OpenStudy (anonymous):

how about nw ?

ganeshie8 (ganeshie8):

\[\large \gcd\left(3434,~~\frac{10^{6n-6}-1}{99} \right)= 1\] you want to prove this for n even number right ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

@Marki ,it's ur idea?

OpenStudy (anonymous):

means ?

OpenStudy (anonymous):

ur doing it for ur main question?

OpenStudy (anonymous):

well tbh , this prove nothing t the previous question :P but its good observation that we always got distinct primes btw those term

OpenStudy (anonymous):

seems to be related :D

OpenStudy (anonymous):

see , CGD btw them is 1 at each time but there sum is composite which makes me think of this x N+y M=1

OpenStudy (anonymous):

but it sounded more like hmm complicated , if i found something ok for sure i'll share :D tell now try to prove this one only

OpenStudy (freckles):

I wish i could stay longer and try on this problem but I have to leave soon I kinda playing with ugly fraction \[\frac{(10^{n-1}-1)(10^{n+1}-1)(10^{2(n-1)}+10^{n-1}+1)(10^{2(n-1)}-10^{n-1}+1)}{(10-1)(10+1)}\]

OpenStudy (anonymous):

sounds interesting !

OpenStudy (freckles):

oops type-0

OpenStudy (anonymous):

i hope u to get something ! all the luck

OpenStudy (freckles):

\[\frac{(10^{n-1}-1)(10^{n-1}+1)(10^{2(n-1)}+10^{n-1}+1)(10^{2(n-1)}-10^{n-1}+1)}{(10-1)(10+1)} \]

OpenStudy (freckles):

I want to cancel the factors on bottom i'm sure we can somehow factor the \[10^{n-1} \pm 1\]

OpenStudy (freckles):

\[3434=(10^2+1)(3 \cdot 10 +4)\]

OpenStudy (freckles):

if n is ever then n-1 is odd so if I'm not mistaken 10^2+1 will not be a factor of the other number

OpenStudy (freckles):

but i'm just sorta thinking too fast right now because i'm rushed i must go now

OpenStudy (anonymous):

yes u can do

ganeshie8 (ganeshie8):

Here is a fancy method : prime factors of \(\large 3434\) are \( \{2, 7, 101\} \) so it is sufficient if we show that the number \(\large \frac{10^{12k-6}-1}{99}\) can never have one of those factors

OpenStudy (anonymous):

:D

OpenStudy (anonymous):

@ganeshie8 wait...

ganeshie8 (ganeshie8):

okay :)

OpenStudy (anonymous):

i'm right working on that idea...

OpenStudy (anonymous):

i think i've proved why it doesn't have the factor 101

OpenStudy (anonymous):

this is what ganesh trying to do

OpenStudy (anonymous):

what's ur idea about 17?

ganeshie8 (ganeshie8):

my idea works for all factors in general but it uses number theory can i see your method for 101 ? :)

OpenStudy (anonymous):

comon , someone type :-|

OpenStudy (anonymous):

for all factors? i consider each of them and then try to say that number doesn't have those factors.... this is what i've done so far (not complete) all of those numbers can be written like this: \(\frac{ 10^6-1 }{99 } = 10101\) 5 digits \(\frac{ 10^18-1 }{ 99 }=101010....1 \) 17 digits \(\frac{ 10^18-1 }{ 99 }=101010....1 \) 29 digits \(10101 \equiv 1 \ (mod \ 101)\) well,i'm using calculator to know what would 1010...1 (17 digits) be...

OpenStudy (anonymous):

awesome!! it's \( 1010..1 (17 \ digits) \equiv 1 \ (mod \ 101) \) too.

OpenStudy (anonymous):

its 1 for any digit ;)

OpenStudy (anonymous):

so relax this can be proven , see second comment

OpenStudy (anonymous):

yes,i know...now i'm working on 17...what's ur idea?

OpenStudy (anonymous):

hmm i cant reveal mine , since i posted this :P lets see what @ganeshie8 have

OpenStudy (anonymous):

but i would like @ganeshie8 to wait a moment..i think i can solve it...

OpenStudy (anonymous):

Hint :- work on mod properties

OpenStudy (anonymous):

i'm exactly working on that , i havn't studied n th for 2 weeks and i think i should first have a look at my book XD

OpenStudy (anonymous):

hmm...the first one is congruent to 3 (mod 17) and second one is congruent to 1 (mod 17 ) :|

ganeshie8 (ganeshie8):

that looks close to what i have :)

OpenStudy (anonymous):

what about mod 10 :D

OpenStudy (anonymous):

all of them are congruent to 1 @Marki ,does this help?

OpenStudy (anonymous):

ok another help if k=17n then k mod 10=7,1,5,9 only

OpenStudy (anonymous):

we need this case so if 17 is a factor then we need this case 17*n=1 mod 10 also n has to have this property n=3 mod 10

OpenStudy (anonymous):

got it! if 17 is a factor of that then it can be written as 17xn and it's congruent to 1,5,7,9 (mod 10) but when is't congruent to 1 then n is congruent to 3 (mod 10) and it cannot ever be happened! is this ur meant?

OpenStudy (anonymous):

and it's odd so doesn't have the factor 2...

OpenStudy (anonymous):

so the question is solved...@ganeshie8 what was ur method?

OpenStudy (anonymous):

why it cant happen ?

OpenStudy (anonymous):

a=17n n=3t + 10 so a=51t + 170

OpenStudy (anonymous):

will continue tomorrow,i'm very tired...

OpenStudy (anonymous):

sry, a=17n n+3t (it can be written as 3t ) --> a = 51t and \( a \equiv 1 \)(mod 10) and it happens whenever t = 1,11,111,1111,... then a would be: 51,561,5661,56661,...

OpenStudy (anonymous):

correct?

OpenStudy (anonymous):

@ganeshie8 , what do u think?

ganeshie8 (ganeshie8):

Looks good to me!

OpenStudy (anonymous):

thanks ;) but i'm not very happy as this question was not such a hard question that i spend a lot of time to solve it , i hope to be better at number th ..

ganeshie8 (ganeshie8):

here is what i had for proving \(17\) cannot divide \(\large \frac{10^{12x-6}-1}{99}\) : we want to show that \(\dfrac{10^{12x-6}-1}{99} = 17k\) is not solvable in integers rearranging we get \(10^{12x-6} = 1+ (99\cdot 17)k\) which is same as showing below congruence has no solution \[10^{12x-6} \equiv 1 \pmod{99\cdot 17}\tag{1}\] Next working out the order of 10 in modulo 99*17 gives 16 so 16 must divide 12x-6 for the congruence to have a solution : \[12x - 6 \equiv 0 \pmod { 16}\] \[2x - 1 \equiv 0 \pmod { 4}\] \[2x - 1 = 4y\] \[2x - 4y = 1\] NO solutions as \(2 \not | ~1\) and consequently \(17\) cannot divide \(\large \frac{10^{12x-6}-1}{99}\)

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

works fine

OpenStudy (anonymous):

eh last one of them

OpenStudy (anonymous):

no,we showed for 101 and 17 and it obviously is not divisible by 2